Main rotor blade
Work continued this week:
- Windscreen washing system adjusted and both sides functioned.
- Annual service paperwork brought up to date.
- Blade anti-node weights removed, inspected and refitted.
- MODs research and paperwork continued.
- Pitot static instrument research into calibration carried out and spares sourced.
- Carrying out BITE tests of AFCS and ASE systems (on going).
- Investigated blade tracking procedure and equipment.
Work to be continued next week:
- Research SI’s that need to be carried out.
- Mods research to be continued.
- Sort and batch parts when they arrive.
twentytwopl 16 March 2021 Uncategorised